# Tailored Forecasting from Short Time Series via Meta-learning

### v1

@@ IV.3.1 The logistic and Gauss iterated maps
For entries with multiple values, those values pertain, from left to right, to our experiments with the logistic map only (Section II.3), the logistic and Gauss iterated maps (Section II.4), and the Lorenz-63 equations (Section II.5).
IV.3.2 The Lorenz-63 equations
To generate the library and test signals, we integrate Eq. 6, using a fourth-order Runge-Kutta scheme with fixed time-step $\Delta t=0.01$, with different values of $v_{1}$ and $\omega_{t}$ and starting from randomly chosen initial conditions. We hold the Lorenz parameters ${v_{2}=28}$ and ${v_{3}=8/3}$ fixed. To ensure that the full duration of each signal lies on its corresponding system’s attractor and contains no transient behavior, we discard the first $1000$ points of each generated trajectory. Except where otherwise indicated, the long library signals consist of ${N_{train}=6000}$ sequential data-points. We use the ${N_{trans}=1000}$ points of each signal as a transient to synchronize the internal state of the forecaster reservoir to the signal and fit the forecaster’s output layer to the remaining ${N_{fit}=5000}$ points. We choose the parameters $v_{1}$ and $\omega_{t}$ of the library members randomly from the uniform distributions ${\omega_{t}\in U[0.75,1.25]}$ and ${v_{1}\in U[7.5,12.5]}$. We choose the test systems’ parameters to form a rectangular grid spanning the region defined by ${0.7\leq\omega_{t}\leq 1.3}$ and ${7\leq v_{1}\leq 13}$. The resolution of this grid is ${30\times 30}$ in Fig. 6 and ${25\times 25}$ for the other figures. We forecast ${N_{for}=3000}$ data-points beyond the end of each test signal. The reservoir hyperparameters (Table 1) for our experiments with the Lorenz-63 Equations were chosen as described for our experiments with the logistic map only and with both the logistic and Gauss iterated maps.
IV.4 Baseline methods for comparison
Except in Section II.5.3, we use the forecaster reservoir hyperparameters given in Table 1 for METAFORS and for all of the baseline methods included in our results. We also discard the same number of data points, $N_{trans}$, before fitting the forecaster RC’s trainable parameters for all methods except Training on the Test Signal. Below we provide additional details for three baseline approaches that are not fully specified in the main text.
IV.4.1 Training on the test signal
We train the forecaster reservoir separately on each short test signal. The reservoir has a zero-vector initial state, ${\bm{r}(0)=\bm{0}}$, at the start of each test signal and we discard the first ${N_{trans}}$ data points before fitting. For our results with the logistic and Gauss iterated maps (Sections II.3 and II.4),
@@ IV.4.1 Training on the test signal
IV.4.2 Multi-task learning
@@
We train the forecaster reservoir on the union of all long library signals. We set the reservoir state to zero at the start of every library signal and discard the first ${N_{trans}}$ data points of each before fitting the forecaster’s output layer.
IV.4.3 Interpolated/extrapolated forecaster
@@
This method relies on knowledge of the dynamical parameters for each of the training and test systems. In our experiments with the logistic and Gauss iterated maps (Sections II.3 and II.4), we implement this method as follows. If the dynamical parameter of the test system is within the range of the the dynamical parameters of the library models, we perform element-wise linear interpolation between the model parameters of the nearest library member with dynamical parameter greater than that of the test system and the nearest library member with dynamical parameter less than that of the test system. If the test system’s dynamical parameter is beyond the range of the library members, we perform element-wise linear extrapolation using the nearest two library members.
@@
In our experiments with the Lorenz-63 equations (Section II.5), we rescale the parameters $\omega_{t}$ and $v_{1}$ associated with each of the library members such that they span a unit interval along both axes. If the dynamical parameters of the test system are within the convex hull of the library, we triangulate the test parameters with respect to those of the library members. Then we perform linear barycentric interpolation of the corresponding forecasters’ trainable model parameters. If the test system is outside the convex hull of the library, we use the forecaster RC of the nearest library member.
Acknowledgements.
We thank Daniel Canaday for helpful conversations, insights, and suggestions. We also acknowledge the University of Maryland supercomputing resources (http://hpcc.umd.edu) made available for conducting the research reported in this paper. This work was supported in part by DARPA under contract W31P4Q-20-C-0077. D.N.’s contributions were also supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 1840340. M.G.’s contributions were also supported by ONR Grant No. N000142212656. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation, the Department of Defense, or the U.S. Government.

### v2

@@ Acknowledgements.
For entries with multiple values, those values pertain, from left to right, to our experiments with the logistic map only, with the logistic and Gauss iterated maps, and with the Lorenz-63 equations.
As shown in  Table 1, the reservoir hyperparameters that we use for our experiments with the logistic map only (Fig. 3) and for our experiments with both the logistic and Gauss iterated maps (Fig. 4) are identical except for the input strengths, $\sigma_{F}$ and $\sigma_{SM}$. We chose the size, mean in-degree, and bias of both networks to have values that typically allow for reasonably accurate forecasting with reservoir computers, and performed no experiment-specific tuning of these values. We chose the leakage rate and spectral radius of the signal mapper such that its memory is sufficiently long that the final reservoir state of the signal mapper, after it has received a test signal as input, is influenced by many, or all, of the data points in the test signal. We chose the input strength of both the forecaster and signal mapper networks such that the average of the product of the input weight matrix and the input data from the library is approximately one, ${\langle B\bm{u}_{i}(t)\rangle_{i,t}=\langle B\bm{L}_{i}(t)\rangle_{i,t}\approx 1}$. We chose the remaining hyperparameters – the regularization strengths, forecaster leakage rate, and forecaster spectral radius – by coarse hand-tuning to allow for good, but not necessarily optimal performance. While more robust hyperparameter tuning may improve performance overall, our priority is to compare the relative performances of simple baseline methods on familiar systems, rather than to obtain highly optimized forecasts.
.2.2 The Lorenz-63 equations
[…]
.3 Baseline methods for comparison
Except in Fig. 5(D), we use the forecaster reservoir hyperparameters given in Table 1 for METAFORS and for all of the baseline methods included in our results. We also discard the same number of data points, $N_{trans}$, before fitting the forecaster RC’s trainable parameters for all methods except Training on the Test Signal. Below we provide additional details for three baseline approaches that are not fully specified in the main text.
.3.1 Training on the test signal
We train the forecaster reservoir separately on each short test signal. The reservoir has a zero-vector initial state, ${\bm{r}(0)=\bm{0}}$, at the start of each test signal and we discard the first ${N_{trans}}$ data points before fitting. For our results with the logistic and Gauss iterated maps,
@@
.3.2 Multi-task learning
@@
We train the forecaster reservoir on the union of all long library signals. We set the reservoir state to zero at the start of every library signal and discard the first ${N_{trans}}$ data points of each before fitting the forecaster’s output layer. This ensures that the reservoir is well synchronized to each long signal for all the data points used for fitting.
.3.3 Interpolated/extrapolated forecaster
@@
This method relies on knowledge of the dynamical parameters for each of the training and test systems. In our experiments with the logistic and Gauss iterated maps (Fig. 3 and Fig. 4), we implement this method as follows. If the dynamical parameter of the test system is within the range of the the dynamical parameters of the library models, we perform element-wise linear interpolation between the model parameters of the nearest library member with dynamical parameter greater than that of the test system and the nearest library member with dynamical parameter less than that of the test system. If the test system’s dynamical parameter is beyond the range of the library members, we perform element-wise linear extrapolation using the nearest two library members.
@@
In our experiments with the Lorenz-63 equations (Fig. 5), we rescale the parameters $\omega_{t}$ and $v_{1}$ associated with each of the library members such that they span a unit interval along both axes. If the dynamical parameters of the test system are within the convex hull of the library, we triangulate the test parameters with respect to those of the library members. Then we perform linear barycentric interpolation of the corresponding forecasters’ trainable model parameters. If the test system is outside the convex hull of the library, we use the forecaster RC of the nearest library member.
Figure S1: METAFORS replicates the logistic map’s dynamics across a large portion of its bifurcation diagram from test signals with unknown dynamical parameters and containing just ${\mathbf{N_{test}=2}}$ data points. We train METAFORS on a library of five trajectories from the logistic map with logistic parameters chosen randomly from ${3.7\leq\mu\leq 3.8}$ (black arrows, F to J). All signals in the library are chaotic; periodic trajectories are excluded from selection. All test signals contain ${N_{test}=2}$ iterations. (A to E) Example short-term forecasts obtained by METAFORS and baseline methods from a test signal with logistic parameters ${\mu_{1}^{*}=3.61}$ and (F to J) bifurcation diagrams constructed by the same methods. Vertical yellow lines indicate values of $\mu$ for which the corresponding forecast leaves the interval $0\leq x\leq 1$ and does not return. In the true bifurcation diagram (F to J), we plot, for each of 500500500 evenly-spaced values of ${2.9\leq\mu\leq 4}$, the final 500500500 iterations of a trajectory of total length 200020002000 iterations starting from a randomly chosen initial condition ${0<x_{0}<1}$. We start each prediction at iteration 100010001000 of the corresponding true trajectory and discard the first 500500500 predicted iterations to ensure that the forecast long-term climate is not obscured on the plot by any initial transient behavior. We plot the subsequent 500500500 predicted iterations.
Figure S2: Even with a more powerful forecaster (with more nodes), our simple multi-task learning approach cannot match METAFORS’ performance. We plot mean valid prediction time, $T_{valid}$, against forecaster size, $N_{F}$, with (A) fully-observed Lorenz systems and (B) partially-observed Lorenz systems ($x_{3}$-only). In (A and B), the test signals have ${N_{test}=200}$ sequential data points (long enough that METAFORS’ cold starting of the forecaster offers no advantage over the other methods). The library consists of the same ${N_{L}=9}$ long signals whose dynamical parameters are marked as black dots in Fig. 5(B) and we calculate mean valid times over a test set of 625625625 time series arranged in a $\times 25$ 25 rectangular grid spanned by ${7\leq v_{1}\leq 13}$ and ${0.7\leq\omega_{t}\leq 1.3}$, as in Section˜.2.2. Error bars denote the standard error of the mean.
Figure S3: The effect of library structure (length and number of long signals) on generalization performance. In Fig. 5(B) we show how the climate-replication performance of METAFORS and a few baseline forecasting methods depends on the relationship between the dynamical parameters of the test and library signals for a sample library containing $N_{L}=9$ long signals of fixed length ${N_{train}=6000}$ (${N_{fit}=5000}$). Here, explore how the structure of the library affects the expected short-term prediction quality over a range of dynamical parameters of the test signals. Namely, we plot mean valid prediction time against (A) the number of library members, $N_{L}$, and (B) the number of data-points in each long library signal used to fit the forecaster’s trainable parameters, $N_{fit}$. In (A), we fix ${N_{fit}=5000}$. In (B), we fix ${N_{L}=9}$. In (A and B), the test signals are of fixed length ${N_{test}=200}$ and the library contains $N_{L}$ long signals with Lorenz parameters chosen randomly from the uniform distributions ${v_{1}\in U[7.5,12.5]}$ and ${\omega_{t}\in U[0.75,1.25]}$. We calculate mean valid times over a test set of 625625625 time series arranged in a $\times 25$ 25 rectangular grid spanned by ${7\leq v_{1}\leq 13}$ and ${0.7\leq\omega_{t}\leq 1.3}$, as in Section˜.2.2. Error bars and shaded regions indicate the standard error of the mean. Independent of the fitting length, $N_{fit}$, we discard a transient of ${N_{trans}=1000}$ data points at the start of each long library signal to train the forecaster RC. Note that the Interpolated/Extrapolated Forecaster requires knowledge of the Lorenz parameters $v_{1}$ and $\omega_{t}$ governing the dynamics of the library and test signals.
Figure S4: METAFORS is useful for cold starting even when generalization is not required. Using a library comprising ${N_{L}=1}$ long time series, this single training signal and all 625625625 test signals are segments of the attractor for the standard Lorenz system, ${\omega_{t}=1}$ and ${v_{1}=10}$, with different initial conditions. Since there is only one library member, the signal mapper learns only a cold-start vector, ${\bm{r}(0)}$, for the forecaster. For all methods, we train the forecaster parameters on the library signal directly. The training and test signals are partially-observed, containing only the $x_{3}$ Lorenz variable. Error bars denote the standard error of the mean. In Backward Extrapolation as a Constant, we extrapolate the test signal backwards from its initial value, ${\bm{s}_{test}(0)}$, as a constant for two-hundred time steps and then synchronize the forecaster to this extrapolated signal
starting from ${\bm{r}(-200\Delta t)=\bm{0}}$. In Training Data Search, we search the training series for the segment, $s_{match}$, that minimizes its root-mean-square distance from the test signal. Denoting by ${\bm{r}_{train}(t_{match})}$ the constructed cold-start vector at the time step of the training signal at which $s_{match}$ begins, we then synchronize the forecaster to the test signal starting from the initial state ${\bm{r}(0)=\bm{r}_{train}(t_{match})}$. Searching the library for the closest match to the test signal offers considerable improvement over the two other elementary approaches that we highlight, but it requires a computationally expensive search each time we wish to make a new prediction. Its performance may also depend on how well the training signal covers its associated attractor. METAFORS, on the other hand, can learn a cold-start vector cheaply for each test signal – no retraining of the signal mapper is required.
Figure S5: Model generalization and cold-starting performance against test signal length in fully-observed Lorenz systems. (A) Using a library comprising $N_{L}=9$ long signals with Lorenz parameters indicated by the black dots in Fig. 5(B) we calculate mean valid times over 625625625 test signals arranged in a $\times 25$ 25 rectangular grid spanning the space defined by ${0.7\leq\omega_{t}\leq 1.3}$ and ${7\leq v_{1}\leq 13}$. For METAFORS, Zero Start (blue), the signal mapper RC learns model parameters but no cold-start vector for the forecaster. For that method and all others except METAFORS, we zero start the forecaster: we synchronize it to the test signal from a zero-vector internal state, ${\bm{r}(0)=\bm{0}}$, and then predict in autonomous/closed-loop mode (Fig. 2B) from the end of the test signal. (B) We use a library comprising ${N_{L}=1}$ long time series. This single training signal and all 625625625 test signals are segments of the attractor for the standard Lorenz system, ${\omega_{t}=1}$ and ${v_{1}=10}$, with different initial conditions. Since there is only one library member, the signal mapper learns only a cold-start vector, ${\bm{r}(0)}$, for the forecaster. For all methods, we train the forecaster parameters on the library signal directly. In (A and B), the training and test signals contain fully-observed Lorenz states and error bars denote the standard error of the mean.
Figure S6: The dependence of generalization performance on Tikhonov regularization strength in fully-observed Lorenz systems. (A) We plot the mean valid time achieved by METAFORS and each of our baseline methods as we vary the regularization strength used to train the forecaster RC. Error bars denote the standard error of the mean. (B) We plot METAFORS’ mean valid prediction time as both the forecaster regularization and the signal mapper regularization vary. In (A) and (B), we calculate mean valid times over a test set of 625625625 time series arranged in a $\times 25$ 25 rectangular grid spanned by ${7\leq v_{1}\leq 13}$ and ${0.7\leq\omega_{t}\leq 1.3}$, as in Section˜.2.2. The ${N_{L}=9}$ long library signals consist of fully-observed Lorenz signals with Lorenz parameters depicted as black dots in Fig. 5(B). The test signals, of $N_{test}=200$ sequential observations, are long enough to synchronize the forecaster’s reservoir state well such that its initialization does not matter. The line labeled METAFORS in plot (A) is a horizontal slice along the line ${\alpha_{SM}=10^{-8}}$ of the heatmap in plot (B). With our reservoir computing implementation of METAFORS (and our chosen reservoir hyperparameters,  Table 1), METAFORS’ generalization benefits from a higher regularization than forecasting fixed dynamics. When training a forecaster RC optimally for prediction of test systems with identical dynamics to the training system, a lower regularization strength ($\alpha_{F}^{Trad}\approx 10^{-13}$) allows for an excellent fit to the training dynamics, and for high valid prediction times. When generalization to new dynamics is our goal, a higher regularization ($\alpha_{F}^{META}\approx 10^{-6}$) prevents over-fitting to the training systems but limits peak performance. Forecaster regularization is thus an important consideration in implementing a METAFORS scheme. The discrepancy that we identify in our Lorenz experiments may not be universal, but demonstrates that regularization strengths that typically offer good performance when generalization is not required may be suboptimal when generalization is necessary.
Figure S7: METAFORS is robust to small amounts of noise. We plot METAFORS’ mean valid prediction time as a function of test signal length and noise-amplitude over 625 test signals with dynamical parameters arranged in a $\times 25$ 25 rectangular grid spanned by ${7\leq v_{1}\leq 13}$ and ${0.7\leq\omega_{t}\leq 1.3}$, as in Section˜.2.2. Error bars denote the standard error of the mean. We add independent and identically distributed Gaussian observational noise with standard deviation, $\sigma_{Noise}$, in amplitude to each test signal before providing it to METAFORS. $\sigma_{Noise}$ is expressed in multiples of the standard deviation in amplitude of all library members, $\sigma_{Lib}$, with both calculated component-wise. Dotted lines: we train METAFORS on the same noiseless library of nine fully-observed long Lorenz signals whose dynamical parameters are depicted as black dots in Fig. 5(B). Dashed lines: we train METAFORS on the same long library signals, but with observational noise of equal amplitude to that of the test signals. In both cases, we measure valid prediction times against noiseless truth signals. Training on signals with noise of equal amplitude to that of the test signals represents the common scenario that the training and test data are both generated by the same process that is noisy or imperfectly measured. When the amplitude of noise in the test signals is low (${\sigma_{Noise}\lesssim 0.01\sigma_{Lib}}$), METAFORS’ performance is strong independent of whether it has been trained on data that also contain noise. If the short test signals contain more significant noise (${\sigma_{Noise}\gtrsim 0.1\sigma_{Lib}}$), METAFORS performance remains robust so long as it has been trained on data with similar levels of noise.